Pecan Theatre Inc. owns and operates movie theaters throughout Florida and Georgia. Pecan Theatre has declared the following annual dividends over a six-year period: Year 1, $24,000; Year 2, $72,000; Year 3, $108,000; Year 4, $132,000; Year 5, $174,000; and Year 6, $216,000. During the entire period ending December 31 of each year, the outstanding stock of the company was composed of 30,000 shares of cumulative, 2% preferred stock, $100 par, and 100,000 shares of common stock, $5 par. Required: 1. Determine the total dividends and the per-share dividends declared on each class of stock for each of the six years. There were no dividends in arrears at the beginning of Year 1. Summarize the data in tabular form. If required, round your answers to two decimal places. If the amount is zero, please enter "0". Preferred Dividends Common Dividends Year Total Dividends Total Per Share Total Per Share Year 1 $ 24,000 $ $ $ $ Year 2 72,000 Year 3 108,000 Year 4 132,000 Year 5 174,000 Year 6 216,000 $ $ 2. Determine the average annual dividend per share for each class of stock for the six-year period. If required, round your answers to two decimal places. Average annual dividend for preferred $ per share Average annual dividend for common $ per share 3. Assuming a market price per share of $156 for the preferred stock and $9 for the common stock, determine the average annual percentage return on initial shareholders' investment, based on the average annual dividend per share for preferred stock and for common stock. Round your answers to two decimal places. Preferred stock % Common stock % Feedback 1. Is the preferred stock cumulative or non-cumulative? How will the preferred stock being cumulative or non-cumulative affect the distribution of dividends? Determine what amount of current dividends the preferred stock should receive per year. Recall the definition of dividends "in arrears". How much in dividends would each share of preferred stock receive each year? How much in dividends would each share of common stock receive each year? 2. Remember you are calculating the average per share for each class of stock, not in total. 3. What would the preferred and common stockholder receive as a return on their initial investment based on the average annual dividend?
Answer #1

1) Dividend distribution :

Preferred dividend = 30000*100*2% = $60000

Preferred Common
Year Total Dividend Total Per share Total Per share
Year 1 24000 24000 0.80 0 0
Year 2 72000 72000 2.40 0 0
Year 3 108000 84000 2.80 24000 0.24
Year 4 132000 60000 2.00 72000 0.72
Year 5 174000 60000 2.00 114000 1.14
Year 6 216000 60000 2.00 156000 1.56
Total 726000 360000 12.00 366000 3.66

2) Average dividend per share of preferred stock = 12/6 = $2 per share

Average dividend per share of common Stock = 3.66/6 = 0.61 per share

3) Preferred dividend = 2/156 = 1.28%

Common Dividend = 0.61/9 = 6.78%